OpenStudy (anonymous):

How do I show this: If f_n(x) = [(x^n(a-bx)^n) /n!] = (b^n / n!)*x^n (pi-x)^n Then show that f_n(pi-x) = f_n(x) I dont even know where to start but its about the irrationality of pi

OpenStudy (anonymous):

what are a and b here?

OpenStudy (anonymous):

they are just integers where pi = a/b

OpenStudy (anonymous):

Is this asking me to place pi in for n and x in for n and then subtract the two?

OpenStudy (anonymous):

I have a question : "pi" is \[\pi\]? Or it is an arbitrary symbol you replace for \[\frac{a}{b} \] ?

OpenStudy (anonymous):

no its pi 3.14

OpenStudy (anonymous):

i think its just a coincidence its also a/b

OpenStudy (anonymous):

why do you have pi=a/b?

OpenStudy (anonymous):

thats what it says in the question "suppose pi = a/b is a rational number with a, b positive integers with no common factor." and then it says the question that i asked

OpenStudy (anonymous):

the point of the question is so i can later in another question prove by contradiction that pi is irrational. So right now we are pretending its rational

OpenStudy (anonymous):

I think you don't need prove by contradiction here because of complex way

OpenStudy (anonymous):

ok i think i got it actually thanks